OpenStudy (peterg2001):

help please will medal For your birthday, you received $325 towards a new laptop that costs $750. You start saving $85 a month. How many months will it take you to save up enough money for the laptop? 3 4 5 6 @ichliebedich @beastieman21 @mover96

OpenStudy (ichliebedich):

so you earn 85 monthly, starting from 325 85x+325 (x=# of months) 85x+325 > or = 750 (you need at least 750 to buy) 85x > or = 425 x=5